19. You want to purchase 50 shares of IBM stock at $80 from your broker. The initial margin is 50%. a. Construct your balanceb.Suppose the stock price falls to $50, construct the balance sheet at that price. (6 points) Liabilities and Equity AssetsC. Calculate the percentage margin at P = $50. If the maintenance margin is 30%, will you receive a margin call? (6 points)d.Suppose you close your position at P = $50. What is your return from this transaction? (6 points)
Answer #1

Solution-)

Net Shares purchased =50

Price =80

Initial margin = 50%

Total investment = 80 * 50 * 50% = 4000 * .5 = 2000

a) Balance Sheet

Assets Liabilities (remains unchanged)

Cash : (2000)

Investment: 2000

b) If the price falls to 50$ , there is no change in the balance sheet since it is a tradeable security it only has its effects in the income statement and not in the balance sheet.

c) Percentage Margin at P=50

=> (80 -50)/80 = .375 = 37.5%

The maintenance margin is 30%, we will not receive the margin call as the trigger price for the same is :

P * (1-Initial Margin)/(1-MM) = 50 * (1-.5)/1-.3 = 50 * .5/.7 = 50 * .714 = 35.71

at Any price of 35.71 or below we will get a margin call and since the price has fallen to 50 only we will not be getting a margin call.

d) If we close the position on 50$ we will be incurring a net loss of

= % Loss * Leverage Ratio = ((80 - 50 ) /80 ) * 1/50% = 30/80 * (1/.5) = .375 * 2 = .75 = 75% loss